OpenStudy (anonymous):

A computer program changed one number into another. The program changes 12 into 9, 20 into 15, and 100 into 75. What number will 60 be changed into? ..

OpenStudy (anonymous):

45, since all the ratios between the numbers is 4/3

OpenStudy (anonymous):

12/9=4/3, 20/15=4/3, 100/75=4/3, thus 60/x also has to be 4/3

OpenStudy (anonymous):

so what would theanswer be

OpenStudy (anonymous):

donuts

OpenStudy (anonymous):

now you have the equation 60/x = 4/3, and all you have to do is simply solve for x. 60/x = 4/3 (multiply by x) 60 = 4x/3 (multiply by 3) 180 = 4x (divide by 4) x = 45

OpenStudy (anonymous):

ok thanks again

OpenStudy (anonymous):

np